# LAMMPS-KOKKOS: Performance Portable Molecular Dynamics Across Exascale Architectures

### v1

@@ 5.1. Single GPU Comparison
[…]
Figure 6. Strong scaling results for LAMMPS on different exascale architectures. The total number of atoms is given by the label below each group of results. In cases of minor atom count discrepancies, the speed was rescaled assuming linear scaling. We see that SNAP, with its faster saturation (4) and higher computational expense delivers the best scaling, while ReaxFF with its lack of saturation plateau struggles to reach even 100 timesteps/s. Overall, we observe excellent strong scaling, even out to 8192 nodes on Frontier and El Capitan.
@@
In 4, we plot performance as a function of number of atoms on a single H100 GPU for the three case studies. “Peak performance” is the point at which short-range molecular dynamics simulations has saturated asymptotic linear scaling in the number of atoms (or run out of HBM). At lower atom counts hardware-induced thread starvation (not enough parallelism) and latency effects, such as kernel launch overheads reduce the achievable performance.
We see that LJ and ReaxFF saturate at a similar point. This trend reflects similar latencies and similar degrees of exposed parallelism. On the other hand, SNAP saturates at much lower atom counts, since the primary compute kernels expose several degrees of parallelism beyond just particle count as described in 4.3.
In 5, we show results from a single GPU/GCD/stack of the three case studies on different hardware for a fixed atom size. While the relative performance of the different architectures does generally follow the generic hardware specifics listed in 1, it is not simply determined by scaling with bandwidth or flop rate. This is not unexpected since molecular dynamics kernels are cache sensitive, have unstructured data access and are not generally compute limited. We suspect that a significant fraction of the difference is explained by the cache size differences. Modern NVIDIA GPUs have significantly larger L1 cache which also can be dynamically leveraged for software managed scratch. We also believe that this difference explains the relatively large jumps between V100 and A100 which exceed the raw performance improvements in bandwidth and flop rate of the GPUs.
@@ 5.2. Exascale System Scalability
In 6 we look at the strong scaling of LAMMPS across all current exascale machines for the three case studies, scaling up to 8192 nodes on OLCF’s Frontier and NNSA’s El Capitan, as well as 2048 nodes on ALCF’s Aurora. We additionally run on 4 GPUs per node on NVIDIA’s Eos DGX Superpod to mimic the largest NVIDIA-based supercomputers. Overall, we see excellent strong scaling performance across all machines and benchmarks, and LAMMPS achieves approximately 1000 timesteps/s for any problem size for LJ and SNAP provided enough nodes are available. SNAP scales particularly well, with its higher computational expense hiding launch latencies and communication, as well as its lower atom count required for GPU saturation (4). Due to its lack of a saturation plateau in 4, ReaxFF shows the poorest scaling, as any node count greater than the smallest required to fit the problem size in memory immediately reduces the efficiency and prevents linear scaling. As a result, no machine is able to exceed 100 timesteps/s for any system size. Finally, we observe that the relative performance of the different machines is consistent with that of the single GPUs shown in 5, with the H100-based Eos machine consistently outperforming the other clusters.
@@ 6. Conclusion
In this paper we discussed the design, optimization strategy, and performance of the LAMMPS KOKKOS package. This package provides a single-source implementation of LAMMPS’s capabilities that is performance-portable across CPUs and multiple vendors’ GPUs via the Kokkos programming model. While that performance portability is critical to be able to support various different architectures without needing to maintain separate code bases, optimization of the implemented algorithms is also essential to achieve good performance. We demonstrated that Kokkos provides the necessary abstractions to leverage the hierarchical nature of parallelism on GPUs, expose hardware capabilities such as software managed caches, and facilitate the optimization of data access patterns. These abstractions facilitated non-trivial algorithmic improvements which give multiplicative performance improvements on top of generational hardware improvements.
In 4 we described several optimizations across three molecular dynamics benchmarks. These optimizations are motivated by GPU design principles: manage memory access patterns, consider arithmetic intensity, maintain code convergence, and expose maximal concurrency.
In 5.1 we showed that, by following these principles, we extract performance across all GPU architectures: multiple generations of NVIDIA GPUs, as well as across GPUs from AMD and Intel. This paper considered a source of performance deviation across architectures of similar capability: cache design. Using cache carveout controls on NVIDIA H100 GPUs, we demonstrated performance drops of 20% to 60% across four top kernels when we matched the L1 cache or shared memory capacity, as appropriate, tothat of AMD MI300A. These deviations are consistent with the performance differences between these GPUs.
In 5.2 we showed that performance on modern supercomputers, including all three US exascale machines, demonstrated excellent scaling. Relative performance was dominated by single GPU performance with network effects subleading, including in the strong scaling regime.
These investigations would not have been possible were it not for the Kokkos single-source performance portable framework and its integration into the LAMMPS software package for molecular dynamics.
@@ Acknowledgements.
We thank Christopher Knight at ALCF for invaluable help with using Aurora. We also thank the many contributors to LAMMPS-KOKKOS, past and present, in particular Ray Shan, Rahul Gayatri, Trung Dac Nguyen, Dan Ibanez, Mitch Murphy, Richard Berger, Nick Curtis, Charlie Sievert, Sikandar Mashaya, Axel Kohlmeyer, and many others listed on the Authors page https://www.lammps.org/authors.html on the LAMMPS website. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology and Engineering Solutions of Sandia, LLC., a wholly owned subsidiary of Honeywell International, Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA-0003525. This research was supported by the U.S. DOE Office of Science-Advanced Scientific Computing Research Program, under Contract No. DE-AC02-06CH11357 and used resources of the Argonne Leadership Computing Facility, a U.S. Department of Energy (DOE) Office of Science user facility at Argonne National Laboratory and is based on research, the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725, and the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility using NERSC award FES-ERCAP0034313.
@@ Appendix A Enabling machine-learning-based interatomic potentials
In addition to advancements in hardware, simulation methodology itself is rapidly evolving. Over the past two decades, and particularly the last few years, one of the main trends in LAMMPS has been a shift away from empirical potentials towards machine-learning-based potentials, such as SNAP (the Spectral Neighbor Analysis Potential) (Thompson et al., 2015) described in 4.3. These methods increase the quantitatively predictive power of the simulations through their foundation of using quantum mechanical calculations to provide atomic forces and energies as training data. At the same time, they are also vastly more computationally expensive than traditional empirical potentials (Plimpton and Thompson, 2012), making the efficient use of modern hardware such as GPUs critical.
LAMMPS and its Kokkos package facilitate different approaches for integrating machine-learning-based potentials. The first and typically most performant strategy is that taken by SNAP, FLARE (Johansson et al., 2022), ACE (Drautz, 2019), and others, where the entire method is reimplemented using the Kokkos library. From a LAMMPS development and compilation perspective, these potentials add no external dependencies and behave as any other potential. The downside to this approach is the associated labor of the manual implementation of the potential, especially its derivatives for force and virial calculation. The vast majority of these labor costs are incurred during initial implementation and have the advantage of being generally robust against other changes in LAMMPS.
@@
LAMMPS supports two integration strategies for potentials implemented using these Python packages. The first is to use the C++ interface to the libraries for those that have one, such the libtorch interface to PyTorch. This requires linking LAMMPS to the corresponding library, having the library be linked to versions of GPU libraries such as CUDA that are compatible with the ones that Kokkos is being linked to, etc., but otherwise these potentials do not require any modifications to LAMMPS itself. This strategy has been adopted by NequIP (Batzner et al., 2022), MACE (Batatia et al., 2022), and others, with Allegro (Musaelian et al., 2023) being the first to combine PyTorch with Kokkos to avoid CPU-GPU data transfer. The second strategy is to embed a Python interpreter in LAMMPS and use it to call the Python libraries, which has the advantage of a potentially simpler installation process. Furthermore, it allows the use of Python packages like JAX which lack a C++ interface altogether. The ML-IAP package in LAMMPS supports this strategy with optional Kokkos acceleration, and is used successfully by HIP-NN (Lubbers et al., 2018).
@@ Appendix B ReaxFF Exascale Preparedness: Integer Overflow
The first case is with the sparse matrix-vector storage format. The number of non-zero values per MPI rank can exceed two billion for sufficiently large local systems. One solution to this is to promote all integer data structures—row offsets, column indices, and row lengths—to 64-bit integers. However, this is needlessly wasteful: the column indices and row lengths are bound by the rank of the matrix, it is only the row offsets that will exceed the range of 32-bit integers on any reasonable system. For this reason the row offsets data structure, which is of length $N_{\text{atoms}}$, where $N_{\text{atoms}}$ is the number of atoms owned by an MPI rank, was promoted to 64-bit integers. The other two integer data structures—most significantly the column indices, which is of length $\approx N_{\text{atoms}}\times N_{\text{neigh}}$, where $N_{\text{neigh}}$ is the maximum number of neighbor atoms—can remain a more space-efficient 32 bit integer View.
@@
The second case was with the bond order and hydrogen bond neighbor table. These were initially implemented as flat 1-d Views where 32-bit integer overflow occurred when indexing into the 1-d View. This was not a trivial refactor because integer offsets into the flat 1-d View were packed into int4 data structures described in the multi-body bonded force section; promoting that to 64-bit integers may have been prescriptive but non-trivial. The more robust solution was to replace the flat 1-d Views with more natural 2-d neighbor tables. Here no index exceeded a 32-bit integer. This still required a code refactor; routines needed to switch from encoding the full 32-bit offset to the neighbor index offset. However this avoided the need to needlessly promote data structures to 64 bit integers.
Appendix C Reproducibility

### v2

@@ 5.1. Single GPU Comparison
Figure 6. Strong scaling results for LAMMPS on different exascale architectures. The total number of atoms is given by the label below each group of results. In cases of minor atom count discrepancies, the speed was rescaled assuming linear scaling. We see that SNAP, with its faster saturation (figure 4) and higher computational expense delivers the best scaling, while ReaxFF with its lack of saturation plateau struggles to reach even 100 timesteps/s. Overall, we observe excellent strong scaling, even out to 8192 nodes on Frontier and El Capitan.
@@
In figure 4, we plot performance as a function of number of atoms on a single H100 GPU for the three case studies. “Peak performance” is the point at which short-range molecular dynamics simulations have saturated asymptotic linear scaling in the number of atoms (or run out of HBM). At lower atom counts, hardware-induced thread starvation (not enough parallelism) and latency effects, such as kernel launch overheads, reduce the achievable performance.
We see that LJ and ReaxFF saturate at a similar point. This trend reflects similar latencies and similar degrees of exposed parallelism. On the other hand, SNAP saturates at much lower atom counts, since the primary compute kernels expose several degrees of parallelism beyond just particle count as described in section 4.3.
In figure 5, we show results from a single GPU/GCD/stack of the three case studies on different hardware for a fixed atom size. While the relative performance of the different architectures does generally follow the generic hardware specifics listed in table 1, it is not simply determined by scaling with bandwidth or flop rate. This is not unexpected since molecular dynamics kernels are cache sensitive, have unstructured data access and are not generally compute limited. We suspect that a significant fraction of the difference is explained by the cache size differences. Modern NVIDIA GPUs have significantly larger L1 cache which also can be dynamically leveraged for software managed scratch. We also believe that this difference explains the relatively large jumps between V100 and A100 which exceed the raw performance improvements in bandwidth and flop rate of the GPUs.
@@ 5.2. Exascale System Scalability
In figure 6 we look at the strong scaling of LAMMPS across all current exascale machines for the three case studies, scaling up to 8192 nodes on OLCF’s Frontier and NNSA’s El Capitan, as well as 2048 nodes on ALCF’s Aurora. We additionally scale up to 2048 nodes of CSCS’s Alps supercomputer and 256 nodes of a 4 GPUs per node on NVIDIA’s Eos DGX Superpod; we intentionally run on four GPUs per Eos node to mimic the configurations of the largest NVIDIA-based supercomputers. [^4] We defer presenting the results from Eos to appendix appendix C for the purpose of clarity and conciseness; the curves for Alps and Eos lie largely on top of each other due to strong similarities between H100 and GH200, as well as the comparable network bandwidths between NDR 400 and Slingshot-11. This is shown in figure 7.
Overall, we see excellent strong scaling performance across all machines and benchmarks, and LAMMPS achieves approximately 1000 timesteps/s for any problem size for LJ and SNAP provided enough nodes are available. SNAP scales particularly well, with its higher computational expense hiding launch latencies and communication, as well as its lower atom count required for GPU saturation (figure 4). Due to its lack of a saturation plateau in figure 4, ReaxFF shows the poorest scaling, as any node count greater than the smallest required to fit the problem size in memory immediately reduces the efficiency and prevents linear scaling. As a result, no machine is able to exceed 100 timesteps/s for any system size. Finally, we observe that the relative performance of the different machines is consistent with that of the single GPUs shown in figure 5.
@@ 6. Conclusion
In this paper we discussed the design, optimization strategy, and performance of the LAMMPS KOKKOS package. This package provides a single-source implementation of LAMMPS’s capabilities that is performance-portable across CPUs and multiple vendors’ GPUs via the Kokkos programming model. While that performance portability is critical to be able to support various different architectures without needing to maintain separate code bases, optimization of the implemented algorithms is also essential to achieve good performance. We demonstrated that Kokkos provides the necessary abstractions to leverage the hierarchical nature of parallelism on GPUs, expose hardware capabilities such as software managed caches, and facilitate the optimization of data access patterns. These abstractions enabled non-trivial algorithmic improvements that give multiplicative performance improvements on top of generational hardware improvements.
In section 4 we described several optimizations across three molecular dynamics benchmarks. These optimizations are motivated by GPU design principles: manage memory access patterns, consider arithmetic intensity, maintain code convergence, and expose maximal concurrency.
In section 5.1 we showed that, by following these principles, we extract performance across all GPU architectures: multiple generations of NVIDIA GPUs, as well as across GPUs from AMD and Intel. This paper considered a source of performance deviation across architectures of similar capability: cache design. Using cache carveout controls on NVIDIA H100 GPUs, we demonstrated performance drops of 20% to 60% across four top kernels when we matched the L1 cache or shared memory capacity, as appropriate, to that of AMD MI300A. These deviations are consistent with the performance differences between these GPUs.
In section 5.2 we showed that performance of the LAMMPS KOKKOS package on modern supercomputers, including three exascale machines, achieves excellent scaling for the three benchmarks. Generally, relative performance was dominated by single GPU performance with network effects subleading. These investigations would not have been possible were it not for the Kokkos single-source performance portable framework and its integration into the LAMMPS software package for molecular dynamics.
@@ Acknowledgements.
We thank Christopher Knight at ALCF for invaluable help with using Aurora. We thank Peter Messmer and Christelle Piechurski at NVIDIA, as well as Maria Grazia Giuffreda and Rocco Meli at CSCS, for their ongoing support with Alps runs at scale. We also thank the many contributors to LAMMPS-KOKKOS, past and present, in particular Ray Shan, Rahul Gayatri, Trung Dac Nguyen, Dan Ibanez, Mitch Murphy, Richard Berger, Nick Curtis, Charlie Sievert, Sikandar Mashaya, Axel Kohlmeyer, and many others listed on the Authors page https://www.lammps.org/authors.html on the LAMMPS website. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology and Engineering Solutions of Sandia, LLC., a wholly owned subsidiary of Honeywell International, Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA-0003525. This research was supported by the U.S. DOE Office of Science-Advanced Scientific Computing Research Program, under Contract No. DE-AC02-06CH11357 and used resources of the Argonne Leadership Computing Facility, a U.S. Department of Energy (DOE) Office of Science user facility at Argonne National Laboratory and is based on research, the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725, and the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility using NERSC award FES-ERCAP0034313.
@@ Appendix A Enabling machine-learning-based interatomic potentials
In addition to advancements in hardware, simulation methodology itself is rapidly evolving. Over the past two decades, and particularly the last few years, one of the main trends in LAMMPS has been a shift away from empirical potentials (such as LJ and ReaxFF) towards machine-learning-based potentials, such as SNAP (the Spectral Neighbor Analysis Potential) (Thompson et al., 2015) described in section 4.3. These methods increase the quantitatively predictive power of the simulations through their foundation of using quantum mechanical calculations to provide atomic forces and energies as training data. At the same time, they are also vastly more computationally expensive than traditional empirical potentials (Plimpton and Thompson, 2012), making the efficient use of modern hardware such as GPUs critical.
LAMMPS and its KOKKOS package facilitate different approaches for integrating machine-learning-based potentials. The first and typically most performant strategy is that taken by SNAP, FLARE (Johansson et al., 2022), ACE (Drautz, 2019), and others, where the entire method is reimplemented using the Kokkos library. From a LAMMPS development and compilation perspective, these potentials add minimal external dependencies and behave as any other potential. The downside to this approach is the associated labor of the manual implementation of the potential, especially its derivatives for force and virial calculation. The vast majority of these labor costs are incurred during initial implementation and have the advantage of being generally robust against other changes in LAMMPS.
@@
LAMMPS supports two integration strategies for potentials implemented using these Python packages. The first is to use the C++ interface to the libraries for those that have one, such as the libtorch interface to PyTorch. This requires linking LAMMPS to the corresponding library, having the library be linked to versions of GPU libraries such as CUDA that are compatible with the ones that Kokkos is being linked to, etc., but otherwise these potentials do not require any modifications to LAMMPS itself. This strategy has been adopted by NequIP (Batzner et al., 2022), MACE (Batatia et al., 2022), and others, with Allegro (Musaelian et al., 2023) being the first to combine PyTorch with Kokkos to avoid CPU-GPU data transfer. The second strategy is to embed a Python interpreter in LAMMPS and use it to call the Python libraries, which has the advantage of a potentially simpler installation process. Furthermore, it allows the use of Python packages like JAX which lack a C++ interface altogether. The ML-IAP package in LAMMPS supports this strategy with optional Kokkos acceleration, and is used successfully by HIP-NN (Lubbers et al., 2018).
@@ Appendix B ReaxFF Exascale Preparedness: Integer Overflow
The first case is with the sparse matrix-vector storage format. The number of non-zero values per MPI rank can exceed two billion for sufficiently large local systems. One solution to this is to promote all integer data structures—row offsets, column indices, and row lengths—to 64-bit integers. However, this is needlessly wasteful: the column indices and row lengths are bound by the rank of the matrix, it is only the row offsets that will exceed the range of 32-bit integers on any reasonable system. For this reason the row offsets data structure, which is of length $N_{\text{atoms}}$, where $N_{\text{atoms}}$ is the number of atoms owned by an MPI rank, was promoted to 64-bit integers. The other two integer data structures—most significantly the column indices, which is of length $\approx N_{\text{atoms}}\times N_{\text{neigh}}$, where $N_{\text{neigh}}$ is the maximum number of neighbor atoms—can remain a more space-efficient 32-bit integer View.
@@
The second case was with the bond order and hydrogen bond neighbor table. These were initially implemented as flat 1-d Views where 32-bit integer overflow occurred when indexing into the 1-d View. This was not a trivial refactor because integer offsets into the flat 1-d View were packed into int4 data structures described in Section 4.2.1; promoting that to 64-bit integers may have been prescriptive but non-trivial. The more robust solution was to replace the flat 1-d Views with more natural 2-d neighbor tables. Here no index exceeded a 32-bit integer. This still required a code refactor; routines needed to switch from encoding the full 32-bit offset to the neighbor index offset. However this avoided the need to needlessly promote data structures to 64-bit integers.
Appendix C Comparisons of Alps and Eos
Figure 7. Strong scaling results for LAMMPS on NVIDIA Eos and CSCS Alps, comparing H100 to GH200 performance across all benchmarks. The data from Alps is identical to the data from figure 6.
In figure 7 we show performance comparisons of Lennard-Jones, ReaxFF, and SNAP between the Alps supercomputer at CSCS (Top500, 2025a) and NVIDIA’s Eos DGX Superpod (Top500, 2025c). We deferred discussion of these machines because of their strong similarity relative to Frontier, El Capitan, and Aurora.
Each node of Alps features four Grace-Hopper superchips (GH200) and Slingshot-11 network in a 1:1 GPU to NIC ratio. Each node of Eos is a DGX H100 featuring eight GPUs per node and NDR 400 for connectivity, again with a 1:1 GPU to NIC ratio. We intentionally only used four GPUs and four NICs per node to match the node configuration of Alps.
Exact hardware differences between H100 and GH200 are enumerated in table 1. The relevant differences between the GPUs is that GH200 has a 20% higher memory bandwidth and capacity than H100; while not noted on the table, it also has 20% higher L2 capacity (50 MiB vs 60 MiB) and a commensurately higher peak L2 throughput. The key similarity is the FP64 performance and unified cache capacities are the same. These similarities and differences lead to different effects on each benchmark, including different behaviors depending on the degree of strong scaling. [^6] Since these benchmarks are fully GPU offloaded, the high CPU-GPU bandwidth on Grace-Hopper relative to PCIe on x86+Hopper has a negligible effect on performance, as does the differences between the Grace CPU on Alps and the Xeon Platinum 8480C 56C on Eos (latency effects aside, as described in the text).
C.1. Lennard-Jones
We see that when we are not deep into the strong-scaling regime, i.e. larger atom-per-GPU problems, Lennard-Jones achieves higher performance on GH200 than H100. This is because the force calculation is L2 throughput limited as opposed to compute bound. This is due to the low compute intensity of the Lennard-Jones potential.
On the other hand, in the deep strong-scaling regime, Eos with H100 outperforms Alps with GH200. This is because of higher launch latencies on GH200 which is more exposed at small per-GPU problem sizes. One approach to amortize these costs is to “reverse” offload some work back to the CPU, inherently removing launch latencies.
This is exposed in LAMMPS via the “-pk kokkos pair/only on” command-line argument, or by manually modifying the suffixes of certain styles so they run on the CPU instead of the GPU (see section 3.3). We did not explore this tuning space as part of these investigations to simplify the comparison, both between H100 and GH200 and separately with MI250X, MI300A, and PVC.
C.2. ReaxFF
For large per-GPU problem sizes, we see a broad similarity in performance between H100 and GH200. This is owing to many of the dominant kernels being compute-limited. One potential exception is the sparse matrix-vector kernel, which in an ideal implementation should be memory bandwidth limited. Profiling indicates that this kernel is limited by latency effects due to the pointer indirection inherent to CSR-like matrix formats. Optimization to remove this limiter is an avenue of further work orthogonal to this paper.
In the deep strong-scaling regime we see Eos outperforming Alps. This is again because of latency effects. ReaxFF has more susceptibility to latency effects because of the several packing and unpacking kernels in the distributed sparse matrix-vector kernel.
C.3. SNAP
The top kernels of the SNAP potential are all either FP64 limited or L1 throughput limited. The performance of each is identical between H100 and GH200. In addition, communications are negligible as a percentage of end-to-end runtime, owing to the high degree of computation in the SNAP force evaluation. For these reasons, performance on Alps and Eos is very close.
Appendix D Reproducibility